Find parametric equations for the curve with the given properties.
The line passing through (6, 5) and (7, 6)
Answer #1

6 t1 7 +1 s 6 17,0 (65) Zet 2 = tit +6 o ostal y = + + + 5 paremetric equation 1 x= t + 6 yos Ostal / y = t+s